McCullough Hospital uses a job-order costing system to assign costs to its patients. Its direct materials include a variety of items such as pharmaceutical drugs, heart valves, artificial hips, and pacemakers. Its direct labor costs (e.g., surgeons, anesthesiologists, radiologists, and nurses) associated with specific surgical procedures and tests are traced to individual patients. All other costs, such as depreciation of medical equipment, insurance, utilities, incidental medical supplies, and the labor costs associated with around-the-clock monitoring of patients are treated as overhead costs.

Historically, McCullough has used one predetermined overhead rate based on the number of patient-days (each night that a patient spends in the hospital counts as one patient-day) to allocate overhead costs to patients. Recently a member of the hospital’s accounting staff has suggested using two predetermined overhead rates (allocated based on the number of patient-days) to improve the accuracy of the costs allocated to patients. The first overhead rate would include all overhead costs within the Intensive Care Unit (ICU) and the second overhead rate would include all Other overhead costs. Information pertaining to the hospital’s estimated number of patient-days, its estimated overhead costs, and two of its patients—Patient A and Patient B—is provided below:

ICU Other Total
Estimated number of patient-days 3,200 16,000 19,200
Estimated fixed overhead cost $ 5,619,200 $ 16,224,000 $ 21,843,200
Estimated variable overhead cost per patient-day $ 260 $ 96
Patient A Patient B
Direct materials $ 5,700 $ 7,400
Direct labor $ 28,000 $ 38,400
Total number of patient-days (including ICU) 21 26
Number of patient-days spent in ICU 0 14

Required:

1. Assuming McCullough uses only one predetermined overhead rate, calculate:

a. The predetermined overhead rate.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.

2. Assuming McCullough calculates two overhead rates as recommended by the staff accountant, calculate:

a. The ICU and Other overhead rates.

b. The total cost, including direct materials, direct labor and applied overhead, assigned to Patient A and Patient B.

(Round "Predetermined overhead rate" to 2 decimal places. Round your intermediate calculations and final answers to the nearest dollar amount.)
Answer #1

1a) Predetermine overhead rate = (3200*260+16000*96+21843200)/19200 = 1261

1b) Total cost assigned

Patient A Patient B
Direct material 5700 7400
Direct labor 28000 38400
Overhead 21*1261 = 26481 26*1261 = 32786
Total cost 60181 78586

2a) ICU overhead rate = (5619200/3200+260) = 2016

Other overhead rate = (16224000/16000+96) = 1110

2b) Total cost assigned

Patient A Patient B
Direct material 5700 7400
Direct labor 28000 38400
ICU overhead 0 14*2016 = 28224
Other Overhead 21*1110 = 23310 12*1110 = 13320
Total cost 57010 87344